A string is wrapped around a uniform solid cylinde
Answer #1

Force balance on block

mg - T = m*a

T = tension in rope

a = acceleration of block

Torque balance around cylinder

Torque = I*alpha = T*r

Icylinder = 0.5*m*r^2

alpha = a/r

0.5*m*r^2*a/r = T*r

T = 0.5*m*a

Using above value

mg - T = m*a

mg - 0.5*m*a = m*a

m*g = 1.5*m*a

a = 2*g/3

Angular acceleration will be

alpha = a/r = 2*g/(3*r)